OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step.(3/2 x)(2/3 y)

Directrix (directrix):

To get the property used in steps which lead to a solution of a problem, a person needs to know the steps that came before. Without those, I don't know the answer. @rick2d2

OpenStudy (anonymous):

\[\frac{ 3 }{2x }*\frac{ 2 }{ 3y }\] cross the 3 on both sides \[\frac{ 1 }{ 2x }*\frac{ 2 }{ y }\] cross the 2 on both sides\[\frac{ 1 }{ x }*\frac{ 1 }{ y }\] take the inverse of it \[\left( \frac{ 1 }{ x }*\frac{ 1 }{ y } \right)^{-1}=x*y\]

OpenStudy (anonymous):

nevermind the last step btw. that one is invalid.

OpenStudy (anonymous):

The end result is \[\frac{ 1 }{ xy}\] assuming i interpret the question right.